OpenStudy (anonymous):

what is the square root of x^4 + 7?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

okay so 4+7 is 11 and the square root of 11 is? what are 2 of the same #s that divide into 11?like the square root of 36 is 6 bcuz 6 times 6 is 36.

OpenStudy (anonymous):

(x^4 + 7)^1/2

OpenStudy (radar):

\[\sqrt{x ^{4}+7}\] or\[\sqrt{x ^{4+7}}\]??

OpenStudy (anonymous):

na I need a specific answer

OpenStudy (anonymous):

it is apparently factorable

OpenStudy (anonymous):

There is no good way to simplify that problem. It is simply \[\sqrt(x^4+7)\] Owlfred, your reply was useless. Helpneededlady, 11 is prime, and that was not this persons question anyway. Jman, is there more to the problem?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

thats it I have to simplify it further

OpenStudy (anonymous):

Find f(√x + 2) if f(x) = x4 + 5

OpenStudy (anonymous):

that is the problem

OpenStudy (anonymous):

only it is x to the fourth

OpenStudy (anonymous):

Well then you already gave us the solution to your problem, and it cannot be simplified further.

OpenStudy (anonymous):

no it has to be simplified

OpenStudy (anonymous):

I did notice some calculation errors, that could be your problem. The answer ends up being \[(\sqrt(x+2))^4+5\]. At this point you can simplify and get a different answer.

OpenStudy (anonymous):

f(√x + 2) if f(x) = x4 + 5 (√x + 2)^4 + 5 = (x+2)^2 +5 = x^2 + 2x +9

OpenStudy (anonymous):

i dont think we r allowed to do that suzi

OpenStudy (anonymous):

why? f(√x + 2) means x = √x + 2 then plug into equation

OpenStudy (anonymous):

suzi, (sqrtx + 2)^4 is not (x + 2)^2

OpenStudy (anonymous):

Guyc is correct, my bad I didn't notice the calculation errors, but her general procedure was correct.

OpenStudy (anonymous):

√x + 2 = (x+2)^1/2 (√x + 2)^4 = ((x+2)^1/2)^4 (x+2)^ 1/2 * 4 (x+2)^2

OpenStudy (anonymous):

it was wrong

OpenStudy (anonymous):

ok up to u, that was my answer and thank you.

OpenStudy (anonymous):

The difficulty we're having is calculating (sqrt x + 2) ^4. Multiplying by hand by brute force I get: x^2 + 8x^(3/2) + 24x + 32x^(1/2) + 16 which is horrible